Platinum Gym has 10,000 gym members out of which 1,500 memberships included Unlimited Fitness Training and use of the tanning salon, and out of which 750 included Unlimited Hydromassage. If the Fitness Training is considered A, the use of the tanning salon is considered B, and the Hydromassage is considered C, then the associate rule for these sales becomes "if A and B are purchased, then C is also purchased."

Given total transactions for C are 3,000. Calculate the lift for this rule.
Answer #1

Total memberships = 10,000 ;Packages of A and B = 1,500 ; Packages of A, B, and C = 750.

The association rule, "If A and B are purchased, then C is also purchased" has a support of 750 out of 1500 sales.

The confidence level = Total support/Total memberships = 750/1,500 = 50%.

The total number of transactions for C is given as 3000 . Benchmark confidence is the number of transactions that include the consequent divided by the total number of transactions.

Therefore the benchmark confidence level = 3,000/10,000 = 30%.

The lift is calculated as confidence / benchmark confidence = 50% / 30% = 1.67.